# Modeling Story Expectations to Understand Engagement: A Generative Framework Using LLMs

### v1

@@ Abstract
Understanding when and why consumers engage with stories is crucial for content creators and platforms. While existing theories suggest that audience beliefs of what is going to happen should play an important role in engagement decisions, empirical work has mostly focused on developing techniques to directly extract features from actual content, rather than capturing forward-looking beliefs, due to the lack of a principled way to model such beliefs in unstructured narrative data. To complement existing feature extraction techniques, this paper introduces a novel framework that leverages large language models to model audience forward-looking beliefs about how stories might unfold. Our method generates multiple potential continuations for each story and extracts features related to expectations, uncertainty, and surprise using established content analysis techniques. Applying our method to over 30,000 book chapters from Wattpad, we demonstrate that our framework complements existing feature engineering techniques by amplifying their marginal explanatory power on average by 31%. The results reveal that different types of engagement—continued reading, commenting, and voting—are driven by distinct combinations of current and anticipated content features. Our framework provides a novel way to study and explore how audience forward-looking beliefs shape their engagement with narrative media, with implications for marketing strategy in content-focused industries.
@@ 4 Data
We scrape free books from Wattpad by focusing on the genres listed on Wattpad’s homepage (e.g., action, adventure, romance). Appendix A provides additional details about the keywords used to scrape the books. For each book, we collect its title and description, when the book was created, the language the book is written in, whether the content is for mature audiences, and writer-provided book tags (e.g., “school”, “drama”, “friendship”). Appendix Table LABEL:tab:book_table shows three book titles and descriptions from our dataset. For each chapter, we collect its text, title, date it was written, and its comment count, vote count, and read count.
We use GPT-4o-mini to summarize the chapters and GPT-3.5-turbo to generate the imagined story continuations.444Specifically, we use GPT-4o-mini-2024-07-18 and GPT-3.5-turbo-0125. We use GPT-4o mini for text summarization because of its good summarization capabilities, its lower cost relative to GPT-3.5-turbo, and because we are not concerned about data leakage for this step. However, for the imagination generation, to alleviate the concern about data leakage (i.e., the engagement metrics and actual story continuations being part of the training data), we only include books that were published after the cutoff date of the training data for GPT-3.5-turbo (September 2021) so all book chapters are published January 2022 and onwards. This ensures the content was not used to train the GPT model we use to generate the imagined stories. Since our proposed story imagination method relies critically on the quality of the input text, we take several steps to clean the scraped stories, which we detail in Appendix B. After cleaning, we are left with 30,258 chapters across 1,735 books.
@@ 7 Conclusion
Appendix A Appendix: Wattpad Data Collection
@@
Our Wattpad dataset is collected through scraping content based on specific search keywords related to the genres listed on Wattpad’s homepage. The genres listed are: Action, Adventure, ChickLit, Classics, Fanfiction, Fantasy, General Fiction, Historical Fiction, Horror, Humor, Mystery, Non-Fiction, Paranormal, Poetry, Random, Romance, Science Fiction, Short Story, Spiritual, Teen Fiction, Thriller, Vampire, Werewolf.
We use ChatGPT to expand the genre list to include a total of 93 keywords: Romance, Fantasy, Science Fiction, Mystery, Thriller, Adventure, Young Adult, Drama, Horror, Historical Fiction, Paranormal, Fanfiction, Dystopian, Comedy, Action, Supernatural, Chick Lit, LGBTQ+, Teen Fiction, New Adult, Short Story, Urban Fantasy, Fairy Tales, Mythology, Detective, Crime, Dark Fantasy, Steampunk, Time Travel, Romantic Suspense, High School, College, Slice of Life, Coming of Age, Magic, Epic Fantasy, Post-Apocalyptic, Cyberpunk, Space Opera, Alien, Zombies, Vampires, Werewolves, Shifters, Demons, Angels, Ghosts, Witches, Wizards, Fairies, Mermaids, Dragons, Alternate History, Military, Political, Psychological, Legal, Medical, Sports, Music, Road Trip, Family, Friendship, Billionaire, Mafia, Royalty, Historical Romance, Western, Inspirational, Spiritual, Self-Help, Memoir, Biography, True Crime, Detective Fiction, Hard-Boiled, Noir, Police Procedural, Legal Thriller, Techno-Thriller, Espionage, Political Thriller, Medical Thriller, Apocalyptic, Humor, Satire, Parody, Tragedy, Love Triangle, Forbidden Love, Soulmates, Enemies to Lovers, Friends to Lovers, Second Chances, Secret Identity, Amnesia, Revenge, Survival, Quest, Heist, Conspiracy, Cult, Secret Society, Virtual Reality, Artificial Intelligence.
Using these 93 keywords, we collect 144,642 chapters across 5,847 books that were published January 2022 and onward. Table LABEL:tab:book_table shares three book titles from our dataset and their descriptions.
| Genre | Title | Description |
| Romance | The Duchess | Seren Ferguson marries the Duke of Blackfall after becoming pregnant with his child. It is, by all accounts, a loveless marriage. Will their relationship improve, or will she find love elsewhere, in unexpected places, from unexpected people? |
| Fantasy | Ashtherielle: The Male Lead’s Older Sister | She died because of cancer, when she woke up she was shocked to find herself as a baby. Later on, she learned her name was Ashtherielle, a minor character of a novel written by her best friend. She made up her mind and accepted the truth, vowing to live until she’s satisfied. |
| Mystery | The Secret Songs of D.C. Silverenn | Cerise Lenoir, a struggling musician with a shopping addiction, discovers that classical music pieces called the Silverenn Songs may hold clues to a mob boss’ hidden treasure. Despite knowing that dangerous competitors are also hunting for the same prize, her desperation for money drives her to join this lethal treasure hunt. |
Appendix B Appendix: Data Cleaning
We clean the stories at the text-level, the chapter-level, and the book-level.
B.1 Cleaning the Text
We remove non-story related text. Specifically, we do the following:
1.
Remove Special Notes: Sometimes writers directly address readers through notes at the beginning or end of the chapter. We remove them as they are not part of the narrative.
2.
Remove HTML Tags and Attributes: We remove HTML tags and attributes since they are not part of the narrative.
B.2 Cleaning the Chapters
After cleaning the text, we remove chapters that are unrelated to the narrative (e.g., preludes, suggested music playlists), conflict with GPT’s text generation policies, or were rewritten, which restarts the read count and distorts the continue-to-read rate.
1.
Remove Irrelevant Chapters: We remove chapters that are not part of the narrative (e.g., author’s notes, playlist suggestions, lists of characters) using several strategies. First, we remove chapters that are titled “introduction,” “prologue,” or “character” or contain the keywords “author” or “extra.” Second, we remove chapters that have a word count less than 2 standard deviations from the average word count of the book since these chapters are unlikely to be part of the narrative. Third, we prompt GPT to provide a score predicting whether the chapter is part of the story or not. In addition to the chapter text, we provide to GPT the chapter title and URL because they can also be informative. We remove chapters that return a probability score less than 0.25, a threshold based on manual inspection of the chapter text. We use the following prompt to determine whether a chapter is part of the story:
“Please classify the following text and provide a probability score between 0 and 1. The score should be close to 1 if the text is a book chapter, and closer to 0 if the text is a prelude, author’s note, Spotify playlist, prologue, or anything else that is not a book chapter. Your response may consist exclusively of a number between 0 and 1. The chapter title is: ‘{title}’. The chapter url is: ‘{url}’.”
2.
Remove Explicit Content: We remove chapters with explicit content since generating explicit content violates OpenAI’s usage policies. To identify explicit content, we prompt GPT to provide a score predicting whether the the chapter contains explicit content. We remove chapters that return a probability score greater than 0.9 as well as the preceding chapter since the continue-to-read rate will be affected. We use the following prompt to determine whether there is explicit content:
“Evaluate the following text and assist me in determining if it features sexually explicit content. Your response may consist exclusively of a probability score from 0 to 1, where a score close to 1 indicates high confidence that the text contains sexually explicit material.”
3.
Remove Rewritten Chapters: Sometimes authors rewrite chapters, which resets the read count for those chapters. We drop chapters that are rewritten as well as the preceding chapter since the continue-to-read rates will reflect the reset read count.
4.
Remove Recently Written Chapters: The data was scraped on July 24, 2024. We remove chapters created after May 24, 2024 because new chapters are likely to provide noisy measures of engagement.
5.
Remove Final Chapter: We remove the last chapter of each book since they do not have a continue-to-read rate. This also removes books with only one chapter.
6.
Remove Chapters with Zero Read Count: We remove chapters that have a read count of zero because the continue-to-read rate will be infinity.
B.3 Cleaning the Books
Finally, the following steps outline which books are removed from the dataset:
1.
Remove Non-English Books: We focus on books written in English because the features we derive from the author-written and GPT-imagined stories require English text. We calculate the percentage of words which are in English for each chapter. We then characterize each book by the average percentage of English words across all chapters. Books that have an average less than 73% are dropped from the dataset. In addition, we drop chapters that are less than 50% English.
2.
Remove Mature Books: In addition to dropping explicit chapters, we drop books that are tagged as “mature” by the author on Wattpad and books that contain explicit content tags (e.g., “adultromance”). We also remove books that have many chapters flagged as explicit since this would lead to large gaps in the narrative. Specifically, we drop books that have more than 15% of its chapters removed due to being explicit content.
3.
Remove Books Unlikely to Have a Narrative: Books with low average word counts tend to lack a narrative structure. For example, several books are just collections of memes. We remove books with an average word count per chapter of less than 400.
4.
Remove Books with High Word Count: We remove books with an average word count per chapter greater than 10,000 because GPT is limited in the length of its input and because often these chapters are not chapters but instead entire books. For the same reason, we also remove chapters with more than 10,000 words.
5.
Remove Books with More Than 50 Chapters: We remove books with a high number of chapters so that books with more chapters do not receive greater weight in our subsequent analysis.
6.
Remove Books with First Chapter Read Count Less Than 100: We remove books with a low read count for the first chapter.

### v2

@@ Abstract
Understanding when and why consumers engage with stories is crucial for content creators and platforms. While existing theories suggest that audience beliefs of what is going to happen should play an important role in engagement decisions, empirical work has mostly focused on developing techniques to directly extract features from actual content, rather than capturing forward-looking beliefs, due to the lack of a principled way to model such beliefs in unstructured narrative data. To complement existing feature extraction techniques, this paper introduces a novel framework that leverages large language models to model audience forward-looking beliefs about how stories might unfold. Our method generates multiple potential continuations for each story and extracts features related to expectations, uncertainty, and surprise using established content analysis techniques. Applying our method to over 30,000 book chapters, we demonstrate that our framework complements existing feature engineering techniques by amplifying their marginal explanatory power on average by 31%. The results reveal that different types of engagement—continued reading, commenting, and voting—are driven by distinct combinations of current and anticipated content features. Our framework provides a novel way to study and explore how audience forward-looking beliefs shape their engagement with narrative media, with implications for marketing strategy in content-focused industries.
@@ 4 Data
We collected a corpus of publicly accessible free books from Wattpad by focusing on the genres listed on Wattpad’s homepage (e.g., action, adventure, romance). Appendix A provides additional details about the keywords used to collect the books. For each book, we collect its title and description, when the book was created, the language the book is written in, whether the content is for mature audiences, and writer-provided book tags (e.g., “school”, “drama”, “friendship”). For each chapter, we collect its text, title, date it was written, and its comment count, vote count, and read count.
We use GPT-4o-mini to summarize the chapters and GPT-3.5-turbo to generate the imagined story continuations.444Specifically, we use GPT-4o-mini-2024-07-18 and GPT-3.5-turbo-0125. We use GPT-4o mini for text summarization because of its good summarization capabilities, its lower cost relative to GPT-3.5-turbo, and because we are not concerned about data leakage for this step. However, for the imagination generation, to alleviate the concern about data leakage (i.e., the engagement metrics and actual story continuations being part of the training data), we only include books that were published after the cutoff date of the training data for GPT-3.5-turbo (September 2021) so all book chapters are published January 2022 and onwards. This ensures the content was not used to train the GPT model we use to generate the imagined stories. Since our proposed story imagination method relies critically on the quality of the input text, we take several steps to clean the collected stories, which we detail in Appendix LABEL:appendix:wattpad_cleaning. After cleaning, we are left with 30,258 chapters across 1,735 books.
@@ 7 Conclusion
Appendix A Appendix: Data Cleaning
@@
Our Wattpad dataset is collected based on specific search keywords related to the genres listed on Wattpad’s homepage. The genres listed are: Action, Adventure, ChickLit, Classics, Fanfiction, Fantasy, General Fiction, Historical Fiction, Horror, Humor, Mystery, Non-Fiction, Paranormal, Poetry, Random, Romance, Science Fiction, Short Story, Spiritual, Teen Fiction, Thriller, Vampire, Werewolf.
We preprocess the dataset at three levels—text, chapter, and book—to focus on content that best reflects narrative structure and reader engagement. This multi-stage cleaning ensures high-quality inputs for downstream analysis and generative modeling.
At the text level, we remove content elements that are clearly outside the core narrative. This includes author commentary, notes directed at readers, and other meta-text that may appear at the beginning or end of chapters. We also remove any formatting tags (e.g., HTML) and other non-linguistic artifacts that may interfere with language model processing.
Chapters are then filtered to retain only those likely to contain meaningful narrative content. We begin by identifying and excluding chapters that are unlikely to be part of the story—such as playlists, author updates, and character lists—using a combination of heuristics (e.g., titles, word count thresholds) and a language model classifier that assigns a likelihood score based on the text and metadata. Chapters with very low predicted narrative probability are excluded. To comply with content policies and ensure model compatibility, we further exclude chapters predicted to contain explicit material. These predictions are made using a language model that estimates the likelihood of such content. To avoid discontinuities in behavioral signals, we also exclude the chapter that immediately precedes flagged ones. We additionally remove chapters that appear to have been rewritten during the data collection window, which can reset engagement counters and distort metrics such as continuation rates. Chapters written shortly before the data collection date are excluded as well, due to limited exposure time and low engagement signal. We also omit final chapters (which lack follow-up continuation metrics) and chapters with zero read counts, which can skew rate-based calculations.
After chapter-level filtering, we apply further criteria at the book level. First, we restrict the dataset to English-language content, estimating language consistency by measuring the proportion of English words in each chapter and computing a book-level average. Books and chapters that fall below language thresholds are removed. We also exclude books tagged with mature content labels, as well as books with a high proportion of chapters predicted to contain sensitive content. This avoids gaps in narrative flow and ensures compatibility with content generation policies. Books with extremely short average chapter lengths (e.g., meme collections or short-form entries) and excessively long chapters (which may reflect full novels pasted into a single entry) are also removed. To maintain consistency in modeling and avoid overrepresentation, we drop books with more than 50 chapters. Finally, we exclude books whose first chapter shows minimal engagement, as this limits our ability to observe reader behavior.